# Flow Matching at Scale: A Machine Learning Framework for Efficient Large-Size Sampling of Many-Body Systems

### v1

@@ Abstract
We propose a machine learning framework based on Flow Matching to overcome the scaling limitations of Markov Chain Monte Carlo (MCMC) methods. We demonstrate its capability in the 2D XY model, where a single network, trained only on configurations from a small ($32\times 32$) lattice at sparse temperature points, generates reliable samples for a significantly larger system ($128\times 128$) across a continuous temperature range without retraining. The generated configurations show strong agreement with key thermodynamic observables and correctly capture the signatures of the Berezinskii-Kosterlitz-Thouless (BKT) transition. This dual generalization is enabled by the Flow Matching framework, which allows us to learn a continuous, temperature-conditioned mapping. At the same time, the inductive biases of the underlying CNN architecture ensure that the learned local physical rules are scale-invariant. This ”train-small, generate-large” capability establishes a new paradigm for efficiently studying critical phenomena, offering a significant computational advantage for exploring the thermodynamic limit. The method can be directly applied to other classical or quantum many-body systems described by continuous fields on a lattice.
@@ I Introduction
Early successes with models like Normalizing Flows (NFs) demonstrated the potential to eliminate critical slowing down for a single set of system parameters [11, 12]. However, the inefficiency of these initial approaches in exploring parameter space is a significant limitation. Studying critical phenomena requires simulations across various parameters (e.g., temperatures), but training a separate, computationally expensive model for each point is impractical [13]. A key advancement to address this has been the development of conditional models, which learn the system’s behavior across a continuous parameter range from sparse training data, enabling efficient interpolation and extrapolation [13, 14]. These models are often implemented with either architecturally rigid NFs, which can suffer from training instabilities like mode collapse [12, 15], or with Diffusion Models (DM), which, while robust, typically rely on a computationally intensive iterative sampling process [16, 17].
@@
The paper is organized as follows. We begin by introducing the theoretical framework of Flow Matching in Sec. II and our simulation setup in Sec. III. Our main numerical results and analyses, covering the model’s dual generalization, performance, and efficiency, are presented in Secs. IV to VII. We then provide a physical interpretation for these capabilities in Sec. VIII. The paper concludes with a discussion of our findings in Sec. IX and an outlook for future work in Sec. X.
@@ II.2 Loss Function
The loss function is similar to the action of the noninteracting fluid system with integration in the entire condition configuration $\phi$ and the initial Gaussian noise. For the path in Eq. (1), conditional velocity field $\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)$ here is just $(\mathbf{x}_{1}-\mathbf{x}_{0})$ and see the Appendix A for more details.
@@ II.3 Flow of Random Sampling
using a method such as explicit Euler or a higher‐order Runge–Kutta. Then, we can transport the samples from the distribution $p_{0}$ to $p_{1}$. For more training and sampling details, see the Appendix B.
@@ III.1 Parameter-Wide Sampling via a Single Network
The training data, consisting of spin configurations $\mathbf{x}_{1}$, uses angles mapped to the range $0,2\pi)$. The neural network is trained to directly output continuous angular values without an explicit modular arithmetic operation (i.e., no wrapping into $(-\pi,\pi]$ or $0,2\pi)$). This is a valid approach because all relevant physical observables, such as the energy and magnetization components, are calculated using trigonometric functions (e.g., $\cos(\Theta_{i}-\Theta_{j})$ and $\sin(\Theta_{i})$), which are inherently periodic with a period of $2\pi$.
@@ V Result II: Reliable Extrapolation of System Sizes
Figure 8: Finite-size scaling of the BKT transition. The susceptibility peak temperature $T_{\text{peak}}(L)$ is plotted against $(\ln L)^{-2}$ for various lattice sizes (L=16, 20, 24, 28, 32, 40, 48, 64, 80, 96, 128L=16,\;20,\;24,\;28,\;32,\;40,\;48,\;64,\;80,\;96,\;128). All data points were generated from a single Flow Matching model trained exclusively on $32\times 32$ data ($T\in[0.1,2.0]$, $\Delta T=0.01$, 1500 configs/T). The linear trend, when extrapolated to the thermodynamic limit ($L\to\infty$), yields $T_{\text{BKT}}\approx 0.898$, in good agreement with the established value of $\approx 0.8935$ [25]. This result confirms that the model has successfully captured the underlying scale-invariant physics.
A finite-size scaling analysis tests the model’s ability to learn scale-invariant physics. Our primary objective here is to qualitatively demonstrate this emergent capability, rather than to achieve a high-precision estimate of $T_{BKT}$. We acknowledge that spin stiffness is a more rigorous metric for analyzing the BKT transition [25]. However, as discussed in our error analysis (Sec. VI), the model’s predictions for fluctuation-related quantities are inherently less precise, a characteristic compounded by our deliberate use of a limited, low-precision training dataset (1500 configurations per temperature from MCMC). Consequently, the statistical noise in the generated stiffness data, particularly near the critical point, was too significant to permit a reliable scaling analysis. We transparently document this behavior alongside other observables in Appendix C.
In contrast, the location of the susceptibility peak, $T_{\text{peak}}(L)$, provided a more robust and clearly identifiable feature for demonstrating the expected scaling behavior. We therefore adopted this metric for our analysis. Using the single model trained on $32\times 32$ data, we located $T_{\text{peak}}(L)$ for various lattice sizes ($L=16$ to $128$). As depicted in Fig. 8, a plot of $T_{\text{peak}}(L)$ versus $(\ln L)^{-2}$ reveals the expected linear relationship [26]. Extrapolating to the thermodynamic limit ($L\to\infty$) yields an estimated $T_{\text{BKT}}\approx 0.898$, in reasonable agreement with the established value of $\approx 0.8935$ [25]. This result, achieved without high-precision training data or retraining, confirms that the model has successfully learned the size-independent physical principles of the XY model, highlighting its efficiency for studying critical phenomena.
@@ VII Result IV: Computational Efficiency for Large-Scale Studies
The primary computational advantage of our Flow Matching approach stems from its ability to generalize across system sizes, a capability traditional MCMC methods lack. We compare the total time cost required for a typical finite-size scaling analysis to quantify the practical implications of this paradigm shift in workflow. For our FM model, this involves a one-time training cost on a small lattice ($32\times 32$), followed by rapid inference on a series of larger lattices. The MCMC method requires running a separate, full simulation for each lattice size. We use our highly optimized, GPU-accelerated MCMC algorithm (see Appendix F) as the baseline for this comparison. All timing benchmarks were performed on a single workstation; detailed hardware and software specifications are provided in Appendix G. The key metric for this evaluation is the Sampling Cost, defined as the average wall-clock time to generate a single statistically independent sample:
@@ VIII.2 U-Net Structure
Mechanistically, the encoder uses an pixel shuffle (space-to-depth) [29] to reduce spatial resolution without discarding information, mapping, for example, each $2\times 2$ block into four channels (see Fig. 12). This preserves translation equivariance and expands the effective receptive field exponentially with depth [30], while convolutions at coarser resolutions still implement local updates on block variables. The decoder applies the inverse pixel shuffle (depth-to-space) to reconstruct high-resolution outputs (see Fig. 12), and skip connections concatenate encoder features with aligned decoder features. This fusion preserves fine-grained locality while enforcing cross-scale consistency, allowing the network to integrate global context without sacrificing the learning of scale-consistent local rules at each resolution. For more architectural and technical details, see the Appendix D.
@@
To empirically validate that the U-Net’s inductive biases are key to its successful size extrapolation, we conducted a comparative study against a Diffusion Transformer (DiT) model [31], which lacks strong, built-in locality. The U-Net significantly outperformed the DiT, especially in the extrapolation task, despite having substantially fewer parameters. This result supports our claim that architectural alignment with the problem’s local nature is critical. The detailed experimental setup, results, and discussion of this comparison are presented in Appendix E.
@@ VIII.3 Learning Local Physics Instead of Global Statistics
Consequently, applying the trained model to a larger lattice does not extrapolate global statistical properties; it simply deploys the same learned local rules over a broader domain. The model generates valid large-scale configurations because the underlying physics it has internalized is the same at all scales. In essence, the generative model did not learn what a small XY system looks like; it learned what the XY model is at a fundamental, local level, thereby enabling the construction of physically plausible configurations of arbitrary size.
@@ IX.3 Limitations of Flow Matching
A significant limitation highlighted by our analysis is the model’s reduced accuracy in predicting second-moment observables, which are directly related to physical fluctuations. This is evident in the increased MSE for susceptibility as system size grows (see Fig. 9) and, more strikingly, in the statistical noise observed in the extrapolated spin stiffness, as detailed in Appendix C. This behavior is likely rooted in the nature of the regression-based training objective (Eq. 10), which excels at learning the mean of the conditional velocity field but may under-represent its complete variance. This characteristic in fact constrains the direct application of this framework to physical problems where high-precision information about fluctuations, such as calculating critical exponents from specific heat or susceptibility, is paramount. Improving the model’s ability to capture the variance of the target distribution faithfully is therefore a critical challenge for future research. Potential avenues could include exploring alternative loss functions that are more sensitive to higher-order moments of the distribution, or incorporating iterative refinement schemes during inference to better sample the whole configuration space.
@@ X Conclusion and Outlook
In this work, we have introduced a generative model based on Flow Matching that represents a new paradigm for the computational study of many-body systems. Our key finding is that a single network, by learning the scale-invariant, local physical rules of the Hamiltonian rather than the global statistics of a specific system size, can generate high-fidelity configurations across continuous temperatures and, crucially, for significantly larger lattices. This dual-generalization capability fundamentally alters the workflow for studying critical phenomena. It replaces the traditional, brute-force approach of running separate, costly simulations for each system size with a more efficient ”train-small, predict-large” strategy. As our analysis shows, this approach provides a significant net computational advantage for large-scale studies, making robust finite-size scaling analysis more accessible.
The principles underpinning this method’s success—aligning the inductive biases of a CNN architecture with the locality inherent in physical Hamiltonians—suggest its applicability to a broader class of models. The framework is particularly well-suited for systems described by continuous fields on a lattice, opening promising avenues for its application in lattice field theories, such as scalar $\phi^{4}$ theory. Furthermore, its utility could extend to quantum many-body problems, especially those where quantum Monte Carlo methods map the system onto a classical statistical model with continuous variables. This work thus establishes a reliable and computationally superior methodology for exploring the thermodynamic limit and provides a flexible foundation for future explorations in statistical mechanics and beyond.
@@ XI Acknowledgment
Appendix A Details of Loss Function
@@
Appendix B Training and Sampling Strategy
@@
Appendix C More Physics Observables in Size Extrapolation
To provide a comprehensive view of the model’s extrapolation performance, this appendix presents a detailed comparison for several key observables across system sizes L=16, 32, 64,L=16,\;32,\;64, and $128$. The model was trained only on $L=32$ data with 1500 configurations per temperature from MCMC. The setup is identical to Sec. V. The results distinguish between the model’s robust performance on first-moment quantities (energy, vortex density) and its challenges with the second-moment, fluctuation-related quantity (spin stiffness).
@@ Appendix C More Physics Observables in Size Extrapolation
Appendix D Neural Network Architecture
@@
D.1 Core Building Blocks
@@
D.2 Optimization of Neural Network Architecture
@@
Appendix E Comparison between U-Net and DiT Architectures
@@
Appendix F GPU-accelerated Monte Carlo Method
We employ a highly optimized MCMC simulation framework implemented in PyTorch [47]. The full implementation of our algorithm is available on GitHub: (https://github.com/ToelUl/Thermal-Phase-Transition-MC-GPU-Simulation). The algorithm is designed to leverage the massively parallel architecture of modern Graphics Processing Units (GPUs), enabling large-scale simulations with high statistical precision. The core design incorporates three key parallelization strategies: a checkerboard update scheme for intra-lattice parallelism, the simulation of multiple independent Markov chains, and the replica-exchange method (also known as Parallel Tempering) across a range of temperatures.
F.1 Parallel Architecture and Data Layout

### v2

@@ Abstract
We propose a machine learning framework based on Flow Matching to overcome the scaling limitations of Markov Chain Monte Carlo (MCMC) methods. We demonstrate its capability in the 2D XY model, where a single network, trained only on configurations from a small ($32\times 32$) lattice at sparse temperature points, generates reliable samples for a significantly larger system ($128\times 128$) across a continuous temperature range without retraining. The generated configurations show strong agreement with key thermodynamic observables and correctly capture the signatures of the Berezinskii-Kosterlitz-Thouless (BKT) transition. This dual generalization is enabled by the Flow Matching framework, which allows us to learn a continuous, temperature-conditioned mapping. At the same time, the inductive biases of the underlying CNN architecture ensure that the learned local physical rules are scale-invariant. This ”train-small, generate-large” capability offers a powerful and efficient alternative for studying critical phenomena. The method can be directly applied to other classical or quantum many-body systems described by continuous fields on a lattice. Furthermore, this framework can serve as a powerful proposal generator in a hybrid scheme with MCMC, dramatically accelerating high-precision studies of the thermodynamic limit.
@@ I Introduction
Early successes with models like Normalizing Flows (NFs) demonstrated the potential to eliminate critical slowing down for a single set of system parameters [11, 13]. However, the inefficiency of these initial approaches in exploring parameter space is a significant limitation. Studying critical phenomena requires simulations across various parameters (e.g., temperatures), but training a separate, computationally expensive model for each point is impractical [14]. A key advancement to address this has been the development of conditional models, which learn the system’s behavior across a continuous parameter range from sparse training data, enabling efficient interpolation and extrapolation [14, 15]. These models are often implemented with either architecturally rigid NFs, which can suffer from training instabilities like mode collapse [13, 16], or with Diffusion Models (DM) [17, 18, 19], which, while robust, typically rely on a computationally intensive iterative sampling process.
@@
The paper is organized as follows. We begin by introducing the theoretical framework of Flow Matching in Sec. II and our simulation setup in Sec. III. Our main numerical results and analyses, covering the model’s dual generalization, performance, and efficiency, are presented in Secs. IV to VII. We then provide a physical interpretation for these capabilities in Sec. VIII. The paper concludes with a discussion of our findings in Sec. IX and an outlook for future work in Sec. X. Further implementation details, more comprehensive results, and the source code for this study are included in the Appendices.
@@ II.2 Loss Function
The loss function is similar to the action of the noninteracting fluid system with integration in the entire condition configuration $\phi$ and the initial Gaussian noise. For the path in Eq. (1), conditional velocity field $\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)$ here is just $(\mathbf{x}_{1}-\mathbf{x}_{0})$ and see the Appendix B for more details.
@@ II.3 Flow of Random Sampling
using a method such as explicit Euler or a higher‐order Runge–Kutta. Then, we can transport the samples from the distribution $p_{0}$ to $p_{1}$. For more training and sampling details, see the Appendix C.
@@ III.1 Parameter-Wide Sampling via a Single Network
The training data, consisting of spin configurations $\mathbf{x}_{1}$, uses angles mapped to the range $0,2\pi)$. The neural network is trained to directly output continuous angular values without an explicit modular arithmetic operation (i.e., no wrapping into $(-\pi,\pi]$ or $0,2\pi)$). This is a valid approach because all relevant physical observables, such as the energy and magnetization components, are calculated using trigonometric functions (e.g., $\cos(\Theta_{i}-\Theta_{j})$ and $\sin(\Theta_{i})$), which are inherently periodic with a period of $2\pi$. The full implementation of our model and the source code for all simulations are made publicly available to ensure reproducibility (see Appendix A).
@@ V Result II: Reliable Extrapolation of System Sizes
Figure 8: Finite-size scaling of the BKT transition. The susceptibility peak temperature $T_{\text{peak}}(L)$ is plotted against $(\ln L)^{-2}$ for various lattice sizes (L=16, 20, 24, 28, 32, 40, 48, 64, 80, 96, 128L=16,\;20,\;24,\;28,\;32,\;40,\;48,\;64,\;80,\;96,\;128). All data points were generated from a single Flow Matching model trained exclusively on $32\times 32$ data ($T\in[0.1,2.0]$, $\Delta T=0.01$, 1500 configs/T). The linear trend, when extrapolated to the thermodynamic limit ($L\to\infty$), yields $T_{\text{BKT}}\approx 0.895$, in good agreement with the established value of $\approx 0.8935$ [26]. This result confirms that the model has successfully captured the underlying scale-invariant physics.
A finite-size scaling analysis tests the model’s ability to learn scale-invariant physics. Our primary objective here is to qualitatively demonstrate this emergent capability, rather than to achieve a high-precision estimate of $T_{BKT}$. We acknowledge that spin stiffness is a more rigorous metric for analyzing the BKT transition [26]. However, as discussed in our error analysis (Sec. VI), the model’s predictions for fluctuation-related quantities are inherently less precise, a characteristic compounded by our deliberate use of a limited, low-precision training dataset (1500 configurations per temperature from MCMC). Consequently, the statistical noise in the generated stiffness data, particularly near the critical point, was too significant to permit a reliable scaling analysis. We transparently document this behavior alongside other observables in Appendix D.
In contrast, the location of the susceptibility peak, $T_{\text{peak}}(L)$, provided a more robust and clearly identifiable feature for demonstrating the expected scaling behavior. We therefore adopted this metric for our analysis. Using the single model trained on $32\times 32$ data, we located $T_{\text{peak}}(L)$ for various lattice sizes ($L=16$ to $128$). As depicted in Fig. 8, a plot of $T_{\text{peak}}(L)$ versus $(\ln L)^{-2}$ reveals the expected linear relationship [27]. Extrapolating to the thermodynamic limit ($L\to\infty$) yields an estimated $T_{\text{BKT}}\approx 0.895$, in reasonable agreement with the established value of $\approx 0.8935$ [26]. This result, achieved without high-precision training data or retraining, confirms that the model has successfully learned the size-independent physical principles of the XY model, highlighting its efficiency for studying critical phenomena.
@@ VII Result IV: Computational Efficiency for Large-Scale Studies
The primary computational advantage of our Flow Matching approach stems from its ability to generalize across system sizes, a capability traditional MCMC methods lack. We compare the total time cost required for a typical finite-size scaling analysis to quantify the practical implications of this paradigm shift in workflow. For our FM model, this involves a one-time training cost on a small lattice ($32\times 32$), followed by rapid inference on a series of larger lattices. The MCMC method requires running a separate, full simulation for each lattice size. We use our highly optimized, GPU-accelerated MCMC algorithm (see Appendix G) as the baseline for this comparison. All timing benchmarks were performed on a single workstation; detailed hardware and software specifications are provided in Appendix H. The key metric for this evaluation is the Sampling Cost, defined as the average wall-clock time to generate a single statistically independent sample:
@@ VIII.2 U-Net Structure
Mechanistically, the encoder uses an pixel shuffle (space-to-depth) [30] to reduce spatial resolution without discarding information, mapping, for example, each $2\times 2$ block into four channels (see Fig. 12). This preserves translation equivariance and expands the effective receptive field exponentially with depth [31], while convolutions at coarser resolutions still implement local updates on block variables. The decoder applies the inverse pixel shuffle (depth-to-space) to reconstruct high-resolution outputs (see Fig. 12), and skip connections concatenate encoder features with aligned decoder features. This fusion preserves fine-grained locality while enforcing cross-scale consistency, allowing the network to integrate global context without sacrificing the learning of scale-consistent local rules at each resolution. For more architectural and technical details, see the Appendix E.
@@
To empirically validate that the U-Net’s inductive biases are key to its successful size extrapolation, we conducted a comparative study against a Diffusion Transformer (DiT) model [32], which lacks strong, built-in locality. The U-Net significantly outperformed the DiT, especially in the extrapolation task, despite having substantially fewer parameters. This result supports our claim that architectural alignment with the problem’s local nature is critical. The detailed experimental setup, results, and discussion of this comparison are presented in Appendix F.
@@ VIII.3 Learning Local Physics Instead of Global Statistics
Consequently, applying the trained model to a larger lattice does not extrapolate global statistical properties; it simply deploys the same learned local rules over a broader domain. The model generates valid large-scale configurations because the underlying physics it has internalized is the same at all scales. In essence, the generative model did not learn what a small XY system looks like; it learned what the XY model is at a fundamental, local level, thereby enabling the construction of physically plausible configurations of arbitrary size. This success underscores a powerful principle in physics-informed machine learning: the most effective models are often those whose architectural inductive biases are deliberately aligned with the intrinsic locality and symmetries of the physical system being studied, a principle that has proven effective in a growing body of literature at the intersection of machine learning and physics [15, 28, 29, 33, 34, 35, 36].
@@ IX.3 Limitations of Flow Matching
A significant limitation highlighted by our analysis is the model’s reduced accuracy in predicting second-moment observables, which are directly related to physical fluctuations. This is evident in the increased MSE for susceptibility as system size grows (see Fig. 9) and, more strikingly, in the statistical noise observed in the extrapolated spin stiffness, as detailed in Appendix D. This behavior is likely rooted in the nature of the regression-based training objective (Eq. 10), which excels at learning the mean of the conditional velocity field but may under-represent its complete variance. This characteristic, in fact, constrains the direct application of this framework to physical problems where high-precision information about fluctuations, such as calculating critical exponents from specific heat or susceptibility, is paramount. Improving the model’s ability to capture the variance of the target distribution faithfully is therefore a critical challenge for future research. While potential avenues could include exploring alternative loss functions, a more immediate and practical solution is integrating our method into a powerful hybrid workflow. In this paradigm, the configurations generated by the FM model serve as high-quality, decorrelated initial states for a subsequent, brief refinement phase using traditional MCMC. A short MCMC chain can efficiently correct for any residual biases from the generative process, ensuring the final ensemble rigorously converges to the target Boltzmann distribution. Combining initial MCMC seeding, massive-scale FM sampling for interpolation and extrapolation, and a final MCMC refinement preserves the core computational advantage. The Flow Matching stage still acts as a potent accelerator, drastically reducing the simulation time required to obtain a large set of independent samples suitable for precision tasks like accurately determining critical exponents.
@@ X Conclusion and Outlook
In this work, we have introduced a generative model based on Flow Matching that offers a powerful alternative for the computational study of many-body systems. Our key finding is that a single network, by learning the scale-invariant, local physical rules of the Hamiltonian rather than the global statistics of a specific system size, can generate high-fidelity configurations across continuous temperatures and, crucially, for significantly larger lattices. This dual-generalization capability fundamentally alters the workflow for studying critical phenomena. It provides an efficient ”train-small, predict-large” strategy that can supplement or replace the traditional, brute-force approach of running separate, costly simulations for each system size. Our analysis shows that this approach provides a significant net computational advantage for large-scale studies, making robust finite-size scaling analysis more accessible.
The principles underpinning this method’s success—aligning the inductive biases of a CNN architecture with the locality inherent in physical Hamiltonians—suggest its applicability to a broader class of models. The framework is particularly well-suited for systems described by continuous fields on a lattice, opening promising avenues for its application in lattice field theories, such as scalar $\phi^{4}$ theory. Furthermore, its utility could extend to quantum many-body problems, especially those where quantum Monte Carlo methods map the system onto a classical statistical model with continuous variables. While our analysis highlights reduced accuracy for second-moment observables, this limitation opens a promising avenue for a powerful hybrid workflow. The core computational advantage is preserved by using the configurations generated by our Flow Matching model as high-quality, decorrelated initial states for a brief MCMC refinement. This synergistic approach, where FM acts as a potent accelerator and MCMC ensures rigorous convergence, paves the way for achieving unprecedented efficiency in high-precision tasks, such as accurately determining critical exponents. This establishes our methodology as an efficient sampling tool and a foundational component for next-generation, hybrid simulation strategies in statistical mechanics and beyond.
@@ XI Acknowledgment
Appendix A Code Availability
The source code for this work’s models, simulations, and analysis is publicly available. The code can be accessed from the GitHub repository at https://github.com/ToelUl/Flow-to-Field. This repository includes the full implementation of the Flow Matching framework, the U-Net architecture, and the GPU-accelerated MCMC simulations. Further details, including hyperparameter settings, random seeds, and instructions for reproducing the figures, are also provided to ensure full reproducibility.
Appendix B Details of Loss Function
@@
Appendix C Training and Sampling Strategy
@@
Appendix D More Physics Observables in Size Extrapolation
To provide a comprehensive view of the model’s extrapolation performance, this appendix presents a detailed comparison for several key observables across system sizes L=16, 32, 64,L=16,\;32,\;64, and $128$. The model was trained only on $L=32$ data with 1500 configurations per temperature from MCMC. The setup is identical to Sec. V. The results distinguish between the model’s robust performance on first-moment quantities (energy in Fig. 14, vortex density in Fig. 15) and its challenges with the second-moment, fluctuation-related quantity (spin stiffness in Fig. 16).
@@ Appendix D More Physics Observables in Size Extrapolation
Appendix E Neural Network Architecture
@@
E.1 Core Building Blocks
@@
E.2 Optimization of Neural Network Architecture
@@
Appendix F Comparison between U-Net and DiT Architectures
@@
Appendix G GPU-accelerated Monte Carlo Method
We employ a highly optimized MCMC simulation framework implemented in PyTorch [50]. While the Wolff [7] or Swendsen–Wang [53] cluster algorithm is known to be more efficient in reducing autocorrelation steps for the 2D XY model, its sequential cluster-building process is not amenable to efficient GPU parallelization. Our choice of the Metropolis-Hastings algorithm with a checkerboard update scheme is motivated by its exceptional performance on massively parallel architectures, making it a powerful and relevant baseline for modern hardware. The full implementation of our algorithm is available on GitHub: (https://github.com/ToelUl/Thermal-Phase-Transition-MC-GPU-Simulation). The algorithm is designed to leverage the massively parallel architecture of modern Graphics Processing Units (GPUs), enabling large-scale simulations with high statistical precision. The core design incorporates three key parallelization strategies: a checkerboard update scheme for intra-lattice parallelism, the simulation of multiple independent Markov chains, and the replica-exchange method (also known as Parallel Tempering [10]) across a range of temperatures.
G.1 Parallel Architecture and Data Layout